# A dataset and model for recognition of audiologically relevant environments for hearing aids: AHEAD-DS and YAMNet+

### v3

@@ Abstract
Scene recognition of audiologically relevant environments is important for hearing aids; however, it is challenging, in part because of the limitations of existing datasets. Datasets often lack public accessibility, completeness, or audiologically relevant labels, hindering systematic comparison of machine learning models. Deploying these models on resource-constrained edge devices presents another challenge. Our solution is two-fold: we leverage several open source datasets to create AHEAD-DS, a dataset designed for scene recognition of audiologically relevant environments, and introduce YAMNet+, a sound recognition model. AHEAD-DS aims to provide a standardised, publicly available dataset with consistent labels relevant to hearing aids, facilitating model comparison. YAMNet+ is designed for deployment on edge devices like smartphones connected to hearing devices, such as hearing aids and wireless earphones with hearing aid functionality; serving as a baseline model for sound-based scene recognition. YAMNet+ achieved a mean average precision of 0.83 and accuracy of 0.93 on the testing set of AHEAD-DS across fourteen categories of audiologically relevant environments. We found that applying transfer learning from the pretrained YAMNet model was essential. We demonstrated real-time sound-based scene recognition capabilities on edge devices by deploying YAMNet+ to an Android smartphone. Even with a Google Pixel 3 (a phone with modest specifications, released in 2018), the model processes audio with approximately 50ms of latency to load the model, and an approximate linear increase of 30ms per 1 second of audio. Our website and code https://github.com/Australian-Future-Hearing-Initiative.
Keywords: scene recognition, sound recognition, audiologically relevant labels, machine hearing, computer audition, AHEAD-DS, YAMNet+
1 Introduction
State of the art hearing devices incorporate scene recognition in order to identify audiological environments and apply appropriate processing (filters, amplification, beam forming, etc) to improve a user’s hearing experience [3].
Several machine learning (ML) models have been proposed for scene recognition [24]. However, there are several limitations: ML models are challenging to port to edge devices [4] and datasets have inconsistent labels (As outlined in Table 1).
@@ 1 Introduction
It is important that models can run on edge devices to preserve privacy by avoiding the transmission of sensitive data and to minimise the cost of data transmission. However, not all models are immediately deployable to devices, such as hearing aids and wireless earphones with hearing aid functionality (henceforth will be collectively referred to as hearing devices), or to portable devices such as smartphones that operate in tandem with hearing devices. Resource constraints and data format compatibility must be overcome for model deployment on the edge [23].
Models are trained and tested using datasets with different scene labels and are thus not directly comparable. These datasets are either not public or incomplete, and general-purpose sound datasets do not have labels that are necessarily relevant to hearing devices [15].
We were motivated to address these problems for several reasons: There is a gap in publicly available standardised datasets for scene recognition of audiologically relevant environments for hearing devices, and an open-source baseline ML model designed for edge devices is currently unavailable. Our ultimate goal is to improve the hearing experience for users of hearing devices. This paper makes the following contributions:
  • A description of a sound dataset named Another HEaring AiD scenes Data Set (AHEAD-DS). This dataset was created for the purpose of scene recognition of audiologically relevant environments for hearing devices, which allows comparison of models using a standardised dataset with consistent labels. The dataset is released under a permissive licence, in a ready-to-use state, divided into training/validation/testing sets and annotated with the ground truth.
  • A description of a ML model called YAMNet+ for sound based scene recognition. The model can act as a baseline model for further development of sound based scene recognition models. The model is released under a permissive licence, including model definitions, training, testing code, and code to convert it into a format deployable to edge devices.
  • We implemented YAMNet+, trained it using AHEAD-DS, and presented benchmark results including precision, recall and inference speed. We have deployed the model to a smartphone to demonstrate it works on an edge device.
This paper is organised as follows. The second section presents an overview of existing datasets and audiologically relevant environment labels. The third section presents an overview of several sound recognition models. The fourth section provides a description of AHEAD-DS and how the dataset was created. The fifth section describes the YAMNet+ sound recognition model. Section six presents the results of training and testing YAMNet+ using AHEAD-DS. Section seven discusses the overall findings of this paper and describes the possible future direction of the research. Section eight draws some conclusions.
2 Scene Recognition Datasets for Hearing Devices
There are several relevant studies (Table 1) involving scene recognition of audiologically relevant environments for hearing devices.
The availability of a standardised dataset helped drive tremendous advancements in computer vision [2]. It is hoped that a dataset for scene recognition for hearing devices will similarly drive advancements. Hence, we sought such a dataset, but noted the following challenges with existing datasets relevant to hearing devices.
  • Datasets may be proprietary, with a restrictive licence and not uploaded to a publicly accessible repository. This was an issue with four studies [33, 8, 11, 39].
  • Data used in the studies were not recorded. Data was processed on-the-fly and raw sound discarded in three studies [14, 18, 5].
  • In one study [15], the dataset available online does not match the description in the published paper.
  • The dataset is a mix of secondary sources, but the specific subset from these sources are unknown. This was an issue for three studies [15, 33, 31].
  • The dataset is a mix of secondary sources but there is ambiguity in the mixing procedure. For example, there may be pseudo random procedures during processing of sound files, which are not clearly explained in the paper. This was an issue with two studies [15, 31].
  • Each study uses datasets with different labels; therefore, they are not easily comparable. This is the case for all studies in Table 1.
@@ 2 Scene Recognition Datasets for Hearing Devices
Additionally, several large general-purpose datasets exist for sound recognition. Datasets such as AudioSet [9], UrbanSound8K [28], ESC50 [27], and TAU scenes [35] provide extensive training material for ML models. However, labels are either too specific or too general to be used directly for the purpose here. For example, in AudioSet recordings can be labelled bus or truck. This level of granular sound recognition is unnecessary: it is only necessary to recognise and respond to motor vehicle noises to improve the hearing experience. However, data can still be used through transfer learning or labels can be remapped to more broad labels.
@@
Conversely, TAU scenes provide location-based recordings which are too general, such as at airports and metro stations. The sound of a motor vehicle will be similarly disruptive irrespective of whether they occur at either location. However, labels could be modified and made more specific, but this would likely require listening to and re-annotating the recordings.
UrbanSound8K and ESC50 provide extensive data. Both datasets contain very specific labels, such as the sounds from specific animals. Like AudioSet, this level of granularity is unnecessary for the purposes here. However, the data is still useful for transfer learning.
Given these issues, we created our own data set using HEAR-DS as a starting point.
Table 1: This table summarises relevant scene labels for hearing devices from several studies. Summarised are the data licence with notes on the data collection and processing procedures.
@@
[15] Cocktail party, Interfering speakers, In traffic, Speech in traffic, In vehicle, Speech in vehicle, Music, Speech in music, Quiet indoors, Speech in qu. Indoors, Reverberant env., Speech in reverb. env. , Wind turbulence, Speech in wind turb. Open source Sound dataset derived from CHiME 5 [1], GTZAN [32], the environment sounds recorded by authors of the study are released publicly. Though the number of samples released is not consistent with the description in the study. The code for mixing speech and environment sounds is missing. Dataset is named Hearing Aid Research Data Set for Acoustic Environment Recognition (HEAR-DS).
[33] Music, Noise, Speech with noise, Speech, Silence Mix of open source and proprietary Sound dataset derived from freesound.org, Demand database [30], CSTR VCTK corpus [37], Million Song Dataset [6], LibriSpeech ASR corpus [26] and YouTube. YouTube source is proprietary.
@@
3 Sound Recognition Machine Learning Models
Three machine learning models relevant for sound recognition are presented in this section, PANN [17], AST [10], and YAMNet [38]. The models in all of these studies were trained, validated, and tested using the AudioSet dataset. The findings are summarised in Table 2. The extensive training dataset makes any of the models suitable for tuning via transfer learning.
While more sound recognition models exist, they are minor variations on the three aforementioned models [40]. Most models make use of frequency domain based features, such as log mel spectrograms. They use either convolution or attention neural network blocks with dense layers. The spectrograms are treated as images and processed using an image processing neural network. Minor variations involve switching spectrograms with log mel spectrograms, or a different configuration of neural network blocks. A minority of models such as PANN use the raw waveform as a feature.
The three models PANN, AST and YAMNet have the following tradeoffs.
  • PANN and AST achieve better mean average precision (mAP) results compared to YAMNet, 0.44 and 0.46 vs 0.31 respectively, using the testing set of AudioSet. However, PANN and AST require over 20 times more parameters compared to YAMNet, approximately 81 and 86 million compared to 3.7 million respectively. This poses a challenge for deploying PANN and AST to resource-constrained edge devices.
  • YAMNet is implemented using the TensorFlow ML framework, while PANN and AST are implemented using the Pytorch ML framework. A robust software pipeline exists to port any TensorFlow model to an edge device, such as an Android phone, using TensorFlow Lite. However, YAMNet’s implementation is not fully open source, while model definitions and weights have been released, training code is missing from the code repositories.
To address these issues, we developed our own model by creating a variation of YAMNet.
Table 2: This table details several machine learning models for sound recognition. The mAP for the AudioSet testing set and number of parameters are those reported by the authors of each study. PANN and AST have higher mAP and use significantly more parameters compared to YAMNet/YAMNet+. This suggests that while PANN and AST have better sound recognition capabilities, they demand significantly more computational resources.
Model Features Architecture Parameters mAP
YAMNet [38] / YAMNet+ Log mel spectrogram Convolution 3,751,369 0.31
PANN [17] Log mel spectrogram, waveform Convolution 81,065,487 0.44
[…]
4 Another HEaring AiD scenes Data Set (AHEAD-DS)
@@
We leverage several open source datasets to create a dataset which we called AHEAD-DS. A summary of the dataset is shown in Table 3. A summary of the sources of the dataset is shown in Table 4. There are a total of 9968 recordings, 10s per recording, across 14 different labels. Each instance is single-channel, 16-bit integers and sampled at 16 kHz.
Despite the limitations outlined earlier, AHEAD-DS is based on HEAR-DS and CHiME 6 Dev [36]. HEAR-DS provides a solid foundation for the following reasons:
  • The recordings of environment sounds have been open sourced with a permissive licence.
  • According to the authors of HEAR-DS, the 14 labels have been chosen from audiologically relevant environments for hearing devices by audiologists and Audifon, a manufacturer of hearing aids.
@@
AHEAD-DS follows the design of HEAR-DS, it employs an identical set of 14 labels. This set comprises six environmental sound categories and an additional six categories representing speech instances within these respective environments. Hearing devices likely require different signal processing strategies to optimise speech intelligibility in each acoustic setting [3]. Therefore, the detection of speech embedded within specific environments represents a step towards creating an improved hearing experience for a hearing devices user.
The main issue with HEAR-DS is that the dataset is not in a fixed, ready-to-use state. This is because the published data is unmixed, and some pseudo-random procedures were used during the production of the HEAR-DS dataset. Since the code has not been published, these procedures cannot be fully reproduced. This causes the following problems.
  • It is not possible to determine the split for training and testing.
  • It is not possible to produce results which can be compared directly against the ones in the HEAR-DS paper.
The advantage of AHEAD-DS is that the dataset is in a ready-to-use state. The code for the mixing procedure is published and all mixing parameters are documented and reproducible. This is important because it removes a variable when comparing ML models. The speech and environment sounds have been released in both mixed and unmixed forms, eliminating any ambiguity regarding the subsets used to derive our new dataset.
@@
The sounds from each environment for AHEAD-DS were taken from HEAR-DS. The authors of HEAR-DS recorded each scene using a dummy head. The head was equipped with a synthetic ear, the GRAS KB1065/1066 Pinnae. An Audiofon microphone was placed in each ear of the dummy head in an in-the-canal position. A Focusrite Scarlett 18i6 sound interface was used for analogue to digital conversion. All recordings were initially captured as 32-bit floating-point values sampled at 48 kHz, and later downsampled to a 16 kHz sampling rate with 16-bit integer quantisation. There were a total of 5929 clips of 10s, from 6 different environments.
@@
One issue raised about AHEAD-DS and HEAR-DS was the constant acoustic signature of the reverberant environment might interfere with the classification of the acoustic signature of the other environments. For sound recognition, if a human can identify the presence of a reverberant environment, an acoustic signature exists for reverberant environments. Deep learning will automatically generate compensating features; these features are issues of concern for manual feature engineering; but not for deep learning [19].
CHiME 6 Dev was used for speech samples in AHEAD-DS instead of CHiME 5, which was used in HEAR-DS. Both contain the same speech recordings, but the start and end times of all recordings have been corrected in CHiME 6 Dev to ensure all recordings are synchronised. CHiME 6 dev is a subset of CHiME 6; it is smaller than the whole dataset but sufficiently long for mixing with environment sounds to provide all the speech in $\langle environment\rangle$ clips and to produce interfering speakers. The interfering speakers clips were derived entirely from CHiME 6 dev.
The authors of CHiME 6 dev recorded the speech in two sessions. Each session in two different home environments consisting of an open space with dining, kitchen, and lounge areas. A total 44 sound files from CHiME 6 Dev were used in AHEAD-DS, 24 recordings from the first environment and 20 in the second. Each session consisted of four people and would begin by having participants take turns to introduce themselves via speaking and then read out a short passage. Once this is complete, participants engage in casual conversation as they prepare and share a meal together.
@@
In CHiME 6 additional speech was recorded using body-worn microphones, but only the recordings from the Kinect were used. The recordings from body-worn microphones lacked acoustic realism as they sounded like clean studio recordings.
The label interfering speakers contains only speech taken from CHiME 6 and is not mixed with environment sounds. It represents a scenario with multiple speakers where the speech is intelligible. The label cocktail party contains multiple speakers but the speech is unintelligible.
Compared to datasets like AudioSet, AHEAD-DS is still relatively small, which is a disadvantage. In later sections, we analyse the implications of having a small dataset (See Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation section).
Table 3: This table summarises the contents of AHEAD-DS. Each recording is 10s, single channel, 16-bit integers and sampled at 16 kHz.
@@
Table 4: AHEAD-DS uses the following data sources from CHiME 6 Dev [36] and HEAR-DS [15].
Source Sound Files Used / Total In Source
CHiME 6 Dev speech 44 / 52
HEAR-DS cocktail party 1334 / 1334
HEAR-DS in traffic 1056 / 1056
HEAR-DS in vehicle 1168 / 1168
HEAR-DS music 2992 / 2992
HEAR-DS quiet indoors 1050 / 1050
HEAR-DS reverb env 443 / 443
HEAR-DS wind turb 878 / 878
4.1 Data Processing Workflow
The data processing workflow is shown in Figure 1.
The sounds sourced from HEAR-DS and CHiME 6 Dev had been encoded in Waveform Audio File Format (WAV) as single-channel, 16-bit integers, and sampled at 16 kHz using several different equipment configurations as described above. AHEAD-DS retained this format.
The levels of CHiME 6 dev and music were high as they were optimised for playback. The signal levels of the other recordings were much lower, and not optimised for playback. The levels were standardised before mixing the clips. The audio files were converted into 32-bit floating-point format during standardisation and mixing, then converted back to 16-bit integers, and saved to disk. For everything except for CHiME 6 Dev and music, the audio was standardised by dividing by the RMS of cocktail party and multiplying by the RMS of CHiME 6 Dev. For music the audio was standardised by dividing by the RMS of music and multiplying by the RMS of CHiME 6 Dev.
These adjustments ensure that the levels of cocktail party and music were approximately similar to conversations in CHiME 6 Dev. Recordings using new hardware configurations must be standardised, this only needs to be done once per device, to ensure recordings of conversations using the new device have similar levels to CHiME 6 Dev. Test results in a later section indicate that this procedure is critical to achieving good performance (See Gain Sensitivity Analysis section).
The first second from each recording in CHiME 6 dev is discarded, as it contains a loud beep to signal the start of the recording. The next 1100s were extracted from each file in 10s intervals. This generated 110 clips per file for a total of 4840 clips of speech.

### v4

# A dataset and model for auditory scene recognition for hearing devices: AHEAD-DS and OpenYAMNet

@@ Abstract
Scene recognition is important for hearing devices, however; this is challenging, in part because of the limitations of existing datasets. Datasets often lack public accessibility, completeness, or audiologically relevant labels, hindering systematic comparison of machine learning models. Deploying such models on resource-constrained edge devices presents another challenge.The proposed solution is two-fold, a repack and refinement of several open source datasets to create AHEAD-DS, a dataset designed for auditory scene recognition for hearing devices, and introduce OpenYAMNet, a sound recognition model. AHEAD-DS aims to provide a standardised, publicly available dataset with consistent labels relevant to hearing aids, facilitating model comparison. OpenYAMNet is designed for deployment on edge devices like smartphones connected to hearing devices, such as hearing aids and wireless earphones with hearing aid functionality, serving as a baseline model for sound-based scene recognition. OpenYAMNet achieved a mean average precision of 0.86 and accuracy of 0.93 on the testing set of AHEAD-DS across fourteen categories relevant to auditory scene recognition. Real-time sound-based scene recognition capabilities were demonstrated on edge devices by deploying OpenYAMNet to an Android smartphone. Even with a 2018 Google Pixel 3, a phone with modest specifications, the model processes audio with approximately 50ms of latency to load the model, and an approximate linear increase of 30ms per 1 second of audio. The project website with links to code, data, and models. https://github.com/Australian-Future-Hearing-Initiative
I Introduction
State of the art hearing devices incorporate scene recognition in order to identify the auditory environment [34]. Scene recognition can be used to trigger distinct hearing device processing strategies [3], such as wind reduction, beamforming, and music mode. Processing strategies improve a hearing device user’s experience by enhancing salient sounds while suppressing undesirable background noise.
@@ I Introduction
Several machine learning (ML) models have been proposed for scene recognition [23]. It is important that models can run locally to preserve privacy, by avoiding the transmission of sensitive data and to minimise the cost of data transmission. Hearing devices like hearing aids, wireless earphones with hearing aid functionality, and smartphones that operate in tandem with hearing devices are edge devices [8]. Porting ML models to edge devices is challenging because such devices are resource constrained [4] and have limited data format compatibility [22].
The challenge of scene recognition is further compounded by the gap in publicly available standardised audio scene recognition datasets. Existing datasets are either in-accessible, incomplete, or not annotated with labels relevant to hearing device listening situations. These limitations prevent the systematic evaluation of ML models for scene recognition.
The motivation for this paper is to address the aforementioned challenges to improve scene recognition. In doing so, allow better application of processing strategies to improve the hearing experience for users of hearing devices. This paper makes the following three contributions. Firstly, a description of a sound dataset named Another HEaring AiD scenes Data Set (AHEAD-DS). This dataset was created for the purpose of scene recognition of the auditory environment, which allows comparison of models using a standardised dataset with consistent labels. The dataset is repacked and refined from several other open source datasets, released under a permissive licence, in a ready-to-use state, divided into training/validation/testing sets and annotated with the ground truth. Secondly, a description of a ML model called OpenYAMNet for sound based scene recognition. The model can act as a baseline model for further development of sound based scene recognition models. The model is released under a permissive licence, including model definitions, training, testing code, and code to convert it into a format deployable to edge devices. Finally, benchmark results of an implementation of the OpenYAMNet ML model, trained using AHEAD-DS, including precision, recall and inference speed. The ML model was deployed to a smartphone to demonstrate the model’s functionality on an edge device.
The remainder of this paper is organised as follows. The remainder of this section presents related work through summaries of existing datasets and ML models. The section that follows presents the methods used to create AHEAD-DS and the OpenYAMNet model. The section after presents test results. The penultimate section discusses the findings of the test results, compares against existing work, and describes the possible future direction of the research. The final section draws some conclusions.
I-A Existing Audio Scene Recognition Datasets for Hearing Devices
Several existing datasets were collected as part of studies relevant to auditory scene recognition, a summary is shown in Table I. The table summarises advantages and issues for individual datasets.
The main challenges of existing audio scene recognition datasets for hearing devices fall into three categories. Public accessibility, completeness, and labelling relevance. Accessibility was a problem for datasets which were immediately discarded after processing, or under a proprietary and restrictive licence. The issue of completeness stems from datasets being a mix of secondary sources, but the mixing procedure and specific subset from secondary sources are unknown.
@@ I-A Existing Audio Scene Recognition Datasets for Hearing Devices
Several large general-purpose datasets exist for sound recognition. Datasets such as AudioSet [9], UrbanSound8K [28], ESC50 [26], and TAU scenes [36] provide extensive training material for ML models. However, the existing labels are either too specific or too general for the objective of auditory scene recognition for hearing devices. For example, AudioSet recordings can be labelled bus or truck. This level of granular sound recognition is unnecessary for a hearing device, it is only necessary to recognise and respond to broad categories like motor vehicle noises to improve the user’s experience. However, this data can still be used through transfer learning, or the labels can be remapped to a more general set.
@@
Conversely, TAU scenes provide location-based recordings which are too general, such as at airports and metro stations. The sound of a motor vehicle will be similarly disruptive irrespective of whether the sound occurs at either location. However, labels could be modified and made more specific, but the modification would likely require listening to and re-annotating the recordings.
Another challenge is the use of different labels across datasets. It is difficult to directly compare ML models trained on each dataset and makes it difficult to combine the datasets into a single, more comprehensive one.
TABLE I: This table summarises relevant datasets for auditory scene recognition for hearing devices.
@@
[16] Cocktail party, Interfering speakers, In traffic, Speech in traffic, In vehicle, Speech in vehicle, Music, Speech in music, Quiet indoors, Speech in quiet indoors, Reverberant env., Speech in reverb. env. , Wind turbulence, Speech in wind turb. Open source Sound dataset derived from CHiME 5 [1], GTZAN [32], the environment sounds recorded by authors of the study are released publicly. Though the number of samples released is not consistent with the description in the study. The code for mixing speech and environment sounds is missing. Dataset is named Hearing Aid Research Data Set for Acoustic Environment Recognition (HEAR-DS).
[34] Music, Noise, Speech with noise, Speech, Silence Mix of open source and proprietary Sound dataset derived from freesound.org, Demand database [30], CSTR VCTK corpus [38], Million Song Dataset [2], LibriSpeech ASR corpus [25] and YouTube. YouTube source is proprietary. The data split and any preprocessing parameters are unknown.
@@
[36] Airport, Bus, Metro, Metro station, Park, Public square, Shopping mall, Street pedestrian, Street traffic, Tram Open source The TAU urban audio visual scenes datasets were recorded in several European cities.”
I-B Existing Sound Recognition Models
Several existing sound recognition models exist, the three presented in this section are Yet Another Mobile Network (YAMNet) [27], Perch [10], and Audio Spectrogram Transformer (AST) [11]. The three representative models were selected as the models provide a good cross-section of common deep learning designs, many other ML models are variations of the three architectures [40].
AST and YAMNet were pre-trained, validated, and tested using the AudioSet dataset, while Perch was pre-trained using XenoCanto [33]. The extensive datasets used for pre-training makes any of the models suitable for fine tuning via transfer learning. AST and YAMNet use log mel spectrograms as features, while Perch uses a per-channel energy normalised mel spectrogram. The spectrograms are represented as two-dimensional inputs and processed using neural networks designed for image recognition.
YAMNet’s implementation is light weight at 3.7 million parameters compared to Perch and AST, which use 7.8 and 86 million parameters respectively. This makes YAMNet a good match for edge computing. However, YAMNet’s implementation is open weight but not fully open source, since only model definitions and weights have been released. The training code for YAMNet is missing from the code repositories, while Perch and AST are fully open source.
II Methods
This section outlines the methodology used to create the AHEAD-DS dataset and the OpenYAMNet model. The procedures for data curation, preprocessing, and labelling for AHEAD-DS are detailed, followed by a description of the OpenYAMNet model architecture, training, and deployment on edge devices.
II-A Another HEaring AiD scenes Data Set (AHEAD-DS)
@@
AHEAD-DS repackages and refines the open source datasets HEAR-DS and CHiME 6 Dev [37]. A summary of the dataset is shown in Table II. There are a total of 9968 recordings, 10s per recording, across 14 different labels. Each instance is single-channel, 16-bit integers and sampled at 16 kHz.
@@
AHEAD-DS follows the design of HEAR-DS and employs an identical set of 14 labels. This set comprises six environmental sound categories and an additional six categories representing speech instances within these respective environments.
@@
The sounds from each environment for AHEAD-DS were taken from HEAR-DS. The authors of HEAR-DS recorded each scene using a dummy head. The head was equipped with a synthetic ear, the GRAS KB1065/1066 Pinnae. An Audifon microphone was placed in each ear of the dummy head in an in-the-canal position. A Focusrite Scarlett 18i6 sound interface was used for analogue to digital conversion. All recordings were initially captured as 32-bit floating-point values sampled at 48 kHz, and later downsampled to a 16 kHz sampling rate with 16-bit integer quantisation. There were a total of 5929 clips of 10s, from 6 different environments.
@@
CHiME 6 Dev was used for speech samples in AHEAD-DS instead of CHiME 5, which was used in HEAR-DS. Both contain the same speech recordings, but the start and end times of all recordings have been corrected in CHiME 6 Dev to ensure all recordings are synchronised. CHiME 6 dev is a subset of CHiME 6, it is smaller than the whole dataset but sufficiently long for mixing with environment sounds to provide all the speech in $\langle environment\rangle$ clips and to produce interfering speakers. The interfering speakers clips were derived entirely from CHiME 6 dev.
The authors of CHiME 6 dev recorded the speech in two sessions. Each session in two different home environments consisting of an open space with dining, kitchen, and lounge areas. A total 44 sound files from CHiME 6 Dev were used in AHEAD-DS, 24 recordings from the first environment and 20 in the second. Each session consisted of four people and would begin by having participants take turns to introduce the participants via speaking and then read out a short passage. Once the introduction is complete, participants engage in casual conversation as the participants prepare and share a meal together.
@@
In CHiME 6 additional speech was recorded using body-worn microphones, but only the recordings from the Kinect were used. The recordings from body-worn microphones lacked acoustic realism as the recordings sounded like clean studio recordings.
The label interfering speakers contains only speech taken from CHiME 6 and is not mixed with environment sounds. The label represents a scenario with multiple speakers where the speech is intelligible. The label cocktail party contains multiple speakers but the speech is unintelligible.
TABLE II: This table summarises the contents of AHEAD-DS. Each recording is 10s, single channel, 16-bit integers and sampled at 16 kHz.
@@
II-B AHEAD-DS Data Processing Workflow
The workflow to create AHEAD-DS is shown in Figure 1. The sounds sourced from HEAR-DS and CHiME 6 Dev had been encoded in Waveform Audio File Format (WAV) as single-channel, 16-bit integers, and sampled at 16 kHz using several different equipment configurations. AHEAD-DS retained this format. For all level adjustments and mixing, the audio files were converted into 32-bit floating-point format during processing. The values were then clipped to the values of $int16_{max}$ and $int16_{min}$ to avoid quantisation errors, then converted back to 16-bit integers, before being saved to disk.
The levels of CHiME 6 dev and music were high as the levels were optimised for playback. The signal levels of the other recordings were much lower, and not optimised for playback. The levels were standardised before mixing the clips. For everything except CHiME 6 Dev and music, the audio was standardised by dividing by the RMS of cocktail party and multiplying by the RMS of CHiME 6 Dev. For music the audio was standardised by dividing by the RMS of music and multiplying by the RMS of CHiME 6 Dev.
These adjustments ensure that the levels of cocktail party and music were approximately similar to conversations in CHiME 6 Dev. Recordings using new hardware configurations must be standardised, this only needs to be done once per device, to ensure recordings of conversations using the new device have similar levels to CHiME 6 Dev.
The first second from each recording in CHiME 6 dev is discarded, as the first second contains a loud beep to signal the start of the recording. The next 1100s were extracted from each file in 10s intervals. This generated 110 clips per file for a total of 4840 clips of speech.